OpenStudy (jazzyfa30):

On the Moon, a falling object falls just 2.65 feet in the first second after being dropped. Each second it falls 5.3 feet farther than in the previous second. How far would an object fall in the first ten seconds after being dropped?

OpenStudy (jazzyfa30):

@bahrom7893

OpenStudy (anonymous):

In the first second it falls 2.65 feet In the second second it falls 2.65 + 5.3 = 7.95 In the third second it falls 7.95 + 5.3 = 13.25 In the fourth second it falls 13.25 + 5.3 = 18.55 In the fifth second it falls 18.55 + 5.3 = 23.85 In the sixth second it falls 23.85 + 5.3 = 29.15 In the seventh second it falls 29.15 + 5.3 = 34.45 In the eighth second it falls 34.45 + 5.3 = 39.75 In the ninth second it falls 39.75 + 5.3 = 45.05 In the tenth second it falls 45.05 + 5.3 = 50.35 Total: 265 feet